OpenStudy (ray10):

use integration by parts: \[\int\limits_{}^{} x \times \ln \left| x \right| dx\]

OpenStudy (bahrom7893):

use u = ln|x|, dv = x

OpenStudy (anonymous):

\[\begin{matrix}u=\ln|x|&dv=x~dx\\ du=\frac{\text{sgn}x}{|x|}&v=\frac{1}{2}x^2\end{matrix}\] where \[\text{sgn}x:=\begin{cases}-1&\text{for }x<0\\0&\text{for }x=0\\1&\text{for }x>0\end{cases}\]

OpenStudy (anonymous):

put lnx=t so x=e^t now differentiate it we get dx=e^t dt now putting it back in the integral we get\[\int\limits e ^{t}t dt\] now integrate by parts

OpenStudy (ray10):

in this case, @bahrom7893 why does \[u = \ln \left| x \right|\] and x=dx?

OpenStudy (anonymous):

$$\int\underbrace{\log|x|}_u\underbrace{x\,\mathrm{d}x}_{\mathrm{d}v}=\underbrace{\frac12 x^2}_v\underbrace{\log|x|}_u-\int \underbrace{\frac12x^2}_v\underbrace{\frac1x\,\mathrm{d}x}_{\mathrm{d}u}=\frac12x^2\log|x|-\frac12\int x\,\mathrm{d}x$$Can you take it from here?

OpenStudy (ray10):

@oldrin.bataku I can take it from there, but I'm wondering how u=ln x, in this case?

OpenStudy (anonymous):

Recall that $$\frac{d}{dx}\log|x|=\frac1x$$ http://en.wikipedia.org/wiki/Natural_logarithm#The_natural_logarithm_in_integration

OpenStudy (anonymous):

@Ray10 we can't integrate things with \(\log\) in them easily at all unless we can substitute them out. Here, our other option is to integrate something comparatively nice (a power of \(x\)) so we chose it instead.

OpenStudy (ray10):

@SithsAndGiggles what is the sgn?

OpenStudy (anonymous):

@Ray10, I wrote the definition in the same comment.

OpenStudy (ray10):

now I understand it more, thanks @SithsAndGiggles

OpenStudy (anonymous):

@Ray10 \(\operatorname{sgn} x\) is the sign of the number; it's \(\operatorname{sgn} 0=0\), \(\operatorname{sgn} x=-1\) for negative \(x\), \(\operatorname{sgn} x=1\) for positive \(x\).

OpenStudy (anonymous):

@SithsAndGiggles is correct but it could be simplified: $$du=\frac{\operatorname{sgn} x}{|x|}=\frac1x$$

OpenStudy (ray10):

so positive \[\frac{ 1 }{ x } = \frac{ sgn x}{ \left| x \right| }\] ?

OpenStudy (anonymous):

forgot the \(dx\)

OpenStudy (anonymous):

@Ray10 they're equal for both positive and negative \(x\); observe;$$\frac{\operatorname{sgn}3}{|3|}=\frac13\\\frac{\operatorname{sgn}(-2)}{|-2|}=\frac{-1}2=\frac1{-2}=-\frac12$$

OpenStudy (ray10):

Taught me something new! @oldrin.bataku appreciate it!

OpenStudy (anonymous):

@Ray10 no problem! glad I could help

OpenStudy (ray10):

and @SithsAndGiggles ! Oh and one question, is there a thanking method on here? like I noticed the medal thing @oldrin.bataku

OpenStudy (anonymous):

Generally you give medals to answers that helped you or that you otherwise liked. You can also leave feedback if you visit their profile.

OpenStudy (ray10):

thank you! :) @oldrin.bataku

OpenStudy (anonymous):

No problem!